Write as the logarithm of a single quantity:
1/4 log(b) 16-2 log(b) 5 + log(b) 7

Answer by venugopalramana(3286)

1/4 log(b) 16-2 log(b) 5 + log(b) 7...USING LOG(X^N)=N*LOG(X)
= {log(b) (16^1/4)}- {log(b) (5^2)} + log(b) 7...USING LOG(XY)=LOG(X)+LOG(Y) AND LOG(X/Y)=LOG(X)-LOG(Y)....AND 2^4=16...AND...5^2=25
= log(b) {2*7/25)=LOG(B)(14/25)